OpenStudy (anonymous):

Find the accumulated value of an investment of $2000 at 8% compounded semiannually for 9 years

OpenStudy (anonymous):

2 856.4925 dollars

OpenStudy (anonymous):

the total amount you'll have including your money

OpenStudy (nikolas):

Using A=P(1+(r/n))^(nt) where A is the total amount, P is the principle (the starting amount), r is the interest rate, n is the number of times per year the interest is compounded, and t is the number of years - A=2000(1+(0.08/2))^(2*9) A=$4051.633 Hmm, this is different from @jygebs' answer, one of us must be doing something wrong... do you know what the correct answer is, @hallbrookerenee ?